OpenStudy (anonymous):

Identify the number as rational or irrational. Explain. 291.87

OpenStudy (saifoo.khan):

Irrational, coz it's a decimal.

OpenStudy (lgbasallote):

i beg to differ...rational numbers are numbers that can be expressed as a quotient of two integers..and correct me if im wrong but i think that can be expressed as 291 87/100 or 29187/100 and i think both are integers yes @saifoo.khan ? ^_^

OpenStudy (saifoo.khan):

1, 1/1 or 1.0 is rational. 1/2, 0.5 is irration.

OpenStudy (saifoo.khan):

irrational*

OpenStudy (anonymous):

It's, in fact, a rational. It can be expressed as \(\left\{\frac{a}{b}\mid a,b,\in\mathbb Z\right\}\).

OpenStudy (lgbasallote):

@saifoo.khan What is a rational number? -Any ordinary number of arithmetic: Any whole number, fraction, mixed number or decimal; together with its negative image. source: http://www.themathpage.com/aprecalc/rational-irrational-numbers.htm

OpenStudy (anonymous):

This is an identity proper only to rationals \(\mathbb Q\) ( http://en.wikipedia.org/wiki/Rational_number ).